1) You titrate 345 mL of 0.125 M propionic acid (Ka = 1.35 × 10–5 ) with 185 mL of 0.165 M KOH. What is the pH of the final solution?

2) What is the pH of a solution formed by titrating 215 mL of 0.45 M nitrous acid (Ka = 7.2 × 10–4 ) with 125.5 mL of 0.855 M NaOH?
Answer #1

1.

Use Henderson Hasselbech equation;

pH = pKa + log [base]/[acid]

pka = -log (ka)

here , pKa = - log(1.35×10^-5) = 4.87

Moles of acid = molarity × volume in litres = 0.125M × 0.345L = 0.0431 moles.

Moles of base (KOH) = 0.165M × 0.185L = 0.0305 moles.

Now put values in initial equation;

pH = 4.87 + log (0.0305)/(0.0431)

pH = 4.87 + log(0.708)

pH = 4.87 + (-0.15) = 4.72

Hence final pH = 4.72